# A Novel Intensity-Based Criterion for Determining Exceptional Point in PT-Symmetry Coupled Array of Optical Waveguides

### v1

@@ 3 Numerical Results
Moreover, the figures distinctly exhibit the characteristic features of discrete diffraction and highlight the impact of boundary reflections. This figure demonstrates the propagation pattern within the Hermitian regime, characterized by the absence of gain or loss ($g=0$). Upon the introduction of gain or loss to the system, and when its magnitude remains below the critical value $g_{c}$, the system stays in the PT-symmetric regime. Figure (3) illustrates that in this regime ($g=0.03$), the propagation pattern is nearly identical to that in the Hermitian regime. However, unlike in the Hermitian state, the total intensity of light in the PT-symmetric phase is not fixed, instead exhibiting oscillations throughout the propagation due to the effects of gain and loss. For better comparison, in all plots related to the propagation pattern, the total intensity is normalized at each propagation distance.
@@
Apart from a factor, the site-to-site correlation is identical for the number state, coherent state, and squeezed state, with the highest values concentrated near the four corners. This result demonstrates that, in the given state and propagation pattern at $z=10$, the probability of detecting two photons in close proximity (either near waveguide 7 or 43) is approximately identical to the probability of detecting two photons at a specific distance — one near waveguide 7 and the other near waveguide 43.
@@ 4 Summary and Conclusion
In addition, the EP was determined using two distinct methods. The results demonstrated consistency between the two approaches. Although computing the eigenvalues of the system is the direct approach for determining the EP, the strong agreement between both methods suggests that the $I_{r}$ criterion, which is based on the propagation pattern, can serve as an effective alternative for identifying the PT-symmetry breaking point. Additionally, the variation of the EP with respect to the number of waveguides was investigated, revealing that the EP value decreases as the number of waveguides in the array increases. In other words, the system enters the broken $\mathcal{PT}$-symmetric phase at a lower gain/loss threshold.

### v2

@@ 3 Numerical Results
Moreover, the figures distinctly exhibit the characteristic features of discrete diffraction [35]. It is important to note that, in the study of optical waveguides array, two common approaches for handling boundaries are zero-boundary conditions and periodic boundary conditions [25, 15, 41, 34]. In this work, we have employed zero-boundary conditions. Since the primary focus of this paper is not on boundary effects, the propagation length and number of waveguides are chosen such that the light does not reach the boundaries during the propagation. Of course, one could investigate a longer propagation length without boundary effects by increasing the number of waveguides. However, the performed study showed that this does not lead to new results and has no significant effect on the overall propagation pattern. Therefore, for the analysis of the propagation pattern and site-to-site correlation, N=50 waveguides and a maximum propagation length of z=10mm are used. Figure 2 demonstrates the propagation pattern within the Hermitian regime, characterized by the absence of gain or loss ($g=0$). Upon the introduction of gain or loss to the system, and when its magnitude remains below the critical value $g_{c}$, the system stays in the PT-symmetric regime. Figure (3) illustrates that in this regime ($g=0.03$), the propagation pattern is nearly identical to that in the Hermitian regime. However, unlike in the Hermitian state, the total intensity of light in the PT-symmetric phase is not fixed, instead exhibiting oscillations throughout the propagation due to the effects of gain and loss. For better comparison, in all plots related to the propagation pattern, the total intensity is normalized at each propagation distance.
@@
It is important to note that Anderson localization, which arises from random disorder, does not occur in our array (with uniform spacing and identical waveguides except for their gain/loss properties) [36, 25, 33]. Furthermore, the existence of defect modes, particularly in the central neutral waveguide of figure 1(b), is highly dependent on the system’s parameters [7]. In our model, with uniform coupling across all waveguides, a detailed investigation into their potential existence is beyond the scope of this study. Moreover, the nature of the input excitation is crucial; in our simulations, only one or two waveguides are excited. This input excites a superposition of the system’s eigenmodes, including both extended and any potential localized modes, rather than selectively exciting localized ones. Consequently, the observed output pattern is a manifestation of the interference and evolution of all modes. The observed delocalized propagation is therefore a direct consequence of the uniform coupling and the specific excitation scheme, and is consistent with the physics of the system.
@@
Apart from a factor, the site-to-site correlation is identical for the number state, coherent state, and squeezed state, with the highest values concentrated near the four corners. This result demonstrates that, in the given state and propagation pattern at $z=10$, the probability of detecting two photons in close proximity (either near waveguide 7 or 43) is approximately identical to the probability of detecting two photons at a specific distance - one near waveguide 7 and the other near waveguide 43.
@@
Interference is the physical core behind how photons distribute and how spatial correlation patterns are formed in a coupled quantum system. In systems featuring gain and loss, interference manifests in more complex forms, which indeed include gain-assisted and loss-assisted interference. These phenomena can be revealed in several ways. Gain can amplify specific pathways for photons, while loss attenuates others. This leads to interference patterns that are unevenly influenced by gain and loss, potentially differing significantly from purely Hermitian or passive systems. Moreover, the presence of gain and loss can induce net photon current in specific directions, which is not typically observed in Hermitian systems. In fact, the observed changes in correlation patterns in the presence of gain and loss are directly attributable to these interferences. In fact, the observed filamentation (patterned structure) in the site-to-site coherence function is a physical phenomenon resulting from interference and the internal gain/loss within the waveguides.
@@ 4 Summary and Conclusion
In addition, the EP was determined using two distinct methods. The results demonstrated consistency between the two approaches. Although computing the eigenvalues of the system is the direct approach for determining the EP, the strong agreement between both methods suggests that the $I_{r}$ criterion, which is based on the propagation pattern, can serve as an effective alternative for identifying the PT-symmetry breaking point. For systems with exceptionally large Hilbert space dimensions, Hamiltonian diagonalization (eigenvalue-based methods [24]) becomes computationally infeasible. This is where the significance of the criterion introduced in this paper becomes apparent, as it can be calculated directly by solving differential equations related to intensity dynamics, without the need for diagonalization. This is particularly critical for complex quantum inputs such as coherent or squeezed states, where the total number of photons is not precisely defined, leading to unmanageable Hilbert space dimensions. Finally, the variation of the EP with respect to the number of waveguides was investigated, revealing that the EP value decreases as the number of waveguides in the array increases. In other words, the system enters the broken $\mathcal{PT}$-symmetric phase at a lower gain/loss threshold.